The contingency table below shows the results of a random sample of 200 state representatives that was conducted to see whether their opinions on a bill are related to their party affiliation. Opinion Approve Disapprove No Opinion 14 18 Republican 42 50 20 24 Democrat Independent 1016 Find the number of degrees of freedon and the critical value x o, to test the claim of independence using a 0.05 A) df 9; x2 16.919 C) df 4; x2-11.143 B) df-5; ?2-11.070 D) df-4.x2-9.488 28) In a X2 test of independence, the null hypothesis is that 28 A) each cell has the same occurrence probability. B) the two variables are dependent. C) the two samples are random. D) the two variables are independent
Answer #1

27)

hre degree of freedom =(row-1)*(column-1) =(3-1)*(3-1)=4

for 4 df and 0.05 level crtiical value =9.488

option D is correct

28)

option D is correct